OpenStudy (anonymous):

y-5=m(x-2) solve for x

OpenStudy (anonymous):

@isael_408 : wlcm to OS.. y-5=m(x-2) y-5=mx-2m y-5+2m=mx so x=(y-5+2m)/m hope you got it

OpenStudy (anonymous):

i do thank you